OpenStudy (anonymous):

Find an equation of the line through (1,5) and parralel to y=3x+5. Write the equation using function notation. f(x)=

OpenStudy (anonymous):

Do I write it out like this: 1+5y=3x+5 correct?

OpenStudy (mathmath333):

no

OpenStudy (anonymous):

y=3x+5(1,5) correct?

OpenStudy (mathmath333):

no

OpenStudy (anonymous):

help :(

OpenStudy (anonymous):

I honestly dont know how to set this equation up

OpenStudy (mathmath333):

slopes of parallel lines are equal \(\large\tt \begin{align} \color{black}{y=3x+5\\ here~~m=3,\\ }\end{align}\) point through which it is parallel is \(\tt \large (1,5)\) therefore the equation of line through which it is parallel is \(\large\tt \begin{align} \color{black}{y-y_1=m(x-x_1)\\ y-5=3(x-1)\\ y-5=3x-3\\ y=3x-3+5\\ y=3x+2\\ }\end{align}\) function is generally considered the y value ,hence \(\large\tt \begin{align} \color{black}{y=3x+2\\~\\ \Large \bf f(x)=3x+2}\end{align}\)

OpenStudy (anonymous):

Can you possibly help me with the second one or tell me if i have the correct answer @mathmath333 ?

OpenStudy (mathmath333):

sry i hve to go, but post it someone can help